# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### 06 Mar 2026 v6

@@ 1 Hilbert’s proof and how to improve it
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof appears in Conrad [3].
Hilbert’s proof is a mathematical gem, but there is still a large room for improvement. The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$, which play a key role in the proof, are uncountable sets. Is it possible to give a proof of the transcendence of $\mathrm{e}$ without substantially using uncountable sets? (For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away. Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible, such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]; they actually proved in this way a much more general result, the Lindemann–Weierstrass theorem. What are the main countable players in their proof which replace the functions $x^{n}\mathrm{e}^{-x}$? The complex sequences
@@
known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
Why do we hear almost nothing about [2]? (In March 2026 it had $1$ citation in Mathematical Reviews.) In the approach of eliminating uncountable sets from proofs, [2] is a remarkable achievement, even if the authors are probably not aware of it (that approach is not mentioned in [2]). However, only a small minority of professional mathematicians, to which this author belongs, care about the distinction of using or not using uncountable sets in proofs; see, for example, the forthcoming book [7].
The proof of Beukers, Bézivin and Robba deserves a new edition, so to say, and we review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$. In Section 3 we present another uncountable-sets-free proof of the transcendence of $\mathrm{e}$ using formal power series. It was found by this author and is based on the natural idea of translating Hilbert’s proof to the language of formal power series.
To conclude the Introduction, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
2 Specializing the proof of Beukers, Bézivin and Robba
@@
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct. For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
@@
The partial fractions are the formal power series
@@
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
@@
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the ordered pairs in the modified recurrence:
@@
From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every number $k\geq k_{0}$ has the required property. $\Box$
@@
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
@@
where $c\geq 0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
@@
In the original proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
@@ 3 The proof by semiformal integration
We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence. For $f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
@@
The family of operations
$\{f(x)\mapsto f(ax)\colon\;a\in\mathbb{R}\}\,$,
which we call constant inner products, consists of formal unary operations. For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
@@
$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
@@
We introduce a family of semiformal unary operations.
@@
We call these operations on $\mathbb{R}[[x]]$ semiformal because coefficients in them arise by limit transitions; in the standard formal operations they are obtained by finite expressions. The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
We show that shifts behave in desirable ways.
@@
In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. Then the coefficients $x^{n}]f(x+c)$ and $x^{n}]g(x+c)$ equal to
$\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,$,}
respectively. Thus
@@
The following semiformal exponential identity is perhaps counter-intuitive; at first we got it wrong.
@@
Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
@@
We use this identity only for a=1a=-1.
We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
Definition 3.7 (Newton integral of fps)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
@@
The third transformation needs some work.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we added
@@
We introduce another family of semiformal functionals on $\mathbb{R}[[x]]$.
Definition 3.11 (improper Newton $\int$ of fps)
@@
We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
@@
We prove Euler’s identity without using uncountable sets.
@@
We are ready to state Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an uncountable-sets-free way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we compute semiformally.
@@
has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most

### 30 Mar 2026 v7

@@ 1 Hilbert’s proof and how to improve it
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$ for $j=0,1,\dots,n$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. Similar proof appears in Conrad [3].
Hilbert’s proof is a mathematical gem, but there is still a large room for improvement. The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$ in Euler’s identity, which are crucial in the proof, are uncountable sets. Is it possible to give a proof of the transcendence of $\mathrm{e}$ that avoids substantial use of uncountable sets? (For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away. Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible; such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]. They proved in this way a much more general result, the Lindemann–Weierstrass theorem. What are the main countable players in their proof replacing the functions $x^{n}\mathrm{e}^{-x}$? The complex sequences
@@
known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal. A nice survey of applications of formal power series is Sambale [6].
The article [2] is little known; in March 2026, it had $1$ citation in Mathematical Reviews. In the perspective of eliminating uncountable sets from proofs, [2] is a remarkable achievement. However, only a small minority of mathematicians, to which this author belongs, care about using or not using uncountable sets in proofs; see, for example, the forthcoming book [8].
The proof of Beukers, Bézivin and Robba deserves a new edition, so to say. We review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$. In Section 3 we present our uncountable-sets-free proof of the transcendence of $\mathrm{e}$. It is based on the translation of Hilbert’s proof to the language of formal power series.
In conclusion, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
2 Specializing the proof of Beukers, Bézivin and Robba in [2]
@@
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct. For $t=0$, the double sum is defined as 0. The numbers $t$, $m_{j}$, and $\alpha_{j}$ come from the factorization
@@
The partial fractions are then the formal power series
@@
(here $x^{n}]f(x)$ denotes the coefficient of $x^{n}$ in $f(x)$) and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
@@
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For pairs $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the pairs in the modified recurrence:
@@
From these three properties, from the implication for the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every number $k\geq k_{0}$ has the required property. $\Box$
@@
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}\ \ (\in\mathbb{C}[[x]])$}
@@
where $c\geq 0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). We know from the proof of Theorem 2.2 that $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
@@
It is clear that the previous proof does not substantially use uncountable sets. In the original proof of the L.–W. theorem in [2], one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
@@ 3 The proof by semiformal integration
In this section, we present an alternative uncountable-sets-free proof of the transcendence of $\mathrm{e}$. It is due to this author, and it is a translation of Hilbert’s proof in the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence. For $f=f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f=[x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
@@
For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations of the derivative $f\mapsto f^{\prime}$ and the primitive $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
@@
$\mathrm{Exp}=\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
@@
We introduce a family of semiformal unary operations on $\mathbb{R}\{x\}$. We call them semiformal because the coefficients of the formal power series involved in these operations are obtained by limit transitions. In the usual formal operations with formal power series, coefficients are obtained by finite expressions.
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations $f(x)\mapsto f(ax)$ and shifts by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
We show that shifts behave as expected.
@@
In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fifth line, we used the binomial theorem. $\Box$
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$. Then
$[x^{n}]f(x+c)=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,[x^{n}]g(x+c)=\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,.$}
Thus
@@
We present a curious semiformal exponential identity.
@@
Proof. Let $f(x)=\mathrm{Exp}(ax)$. Then
@@
Later we use this identity only with a=1a=-1.
We introduce a family of semiformal functionals on $\mathbb{R}\{x\}$.
[…]
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ to be the difference
@@
To prove the third transformation, some work is needed.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, using the change of variable $l=m+1$, we added
@@
We introduce another family of semiformal functionals on $\mathbb{R}\{x\}$.
[…]
@@
We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them in the next bound, which is based on two values of the primitive.
@@
We state and prove Euler’s identity without using uncountable sets.
@@
We are ready to phrase Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are as in Section 1, only now they are defined in an uncountable-sets-free way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, only now we compute semiformally in $\mathbb{R}\{x\}$.
@@
has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15, each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
@@
This translation is more complicated than the $1\frac{1}{2}$ page proof of Hilbert at the beginning, but we eliminated uncountable sets. We think that the semiformal techniques used are of independent interest and will find other applications.